OpenStudy (anonymous):

PLEASE HELP!! in how many orders can 7 chapters be arranged in the index of a textbook if the chapter on probability is to be listed in the middle of the series of chapters???? a. 720 b. 5040 c. 46,656 d. 823,543

OpenStudy (anonymous):

since you have determined one of the positions, this is the same as asking how many ways can you arrange 6 chapters

OpenStudy (anonymous):

is there some kind of formula??

OpenStudy (anonymous):

 do you know how to do those

OpenStudy (anonymous):

NO! not at all! can you show me?

OpenStudy (anonymous):

\[6! \] is the succinct way to write it, but it is the counting principle that gets you the result 6 choices for first slot 5 for second 4 for third 3 for fourth 2 for fifth 1 for sixth counting principle says multiply these together, get \[6\times 5\times 4\times 3\times 2\times 1\]

OpenStudy (anonymous):

720

OpenStudy (anonymous):

thank you so much!!